# Safe Flow Matching: Robot Motion Planning with Control Barrier Functions

### v1

@@ Abstract
Recent advances in generative modeling have led to promising results in robot motion planning, particularly through diffusion and flow-based models that capture complex, multimodal trajectory distributions. However, these methods are typically trained offline and remain limited when faced with unseen environments or dynamic constraints, often lacking explicit mechanisms to ensure safety during deployment. In this work, we propose, Safe Flow Matching (SafeFM), a motion planning approach for trajectory generation that integrates flow matching with safety guarantees. By incorporating the proposed flow matching barrier functions, SafeFM ensures that generated trajectories remain within safe regions throughout the planning horizon, even in the presence of previously unseen obstacles or state-action constraints. Unlike diffusion-based approaches, our method allows for direct, efficient sampling of constraint-satisfying trajectories, making it well-suited for real-time motion planning. We evaluate SafeFM on a diverse set of tasks, including planar robot navigation and 7-DoF manipulation, demonstrating superior safety, generalization, and planning performance compared to state-of-the-art generative planners. Comprehensive resources are available on the project website: https://safeflowmatching.github.io/SafeFM/
@@ 1 Introduction
Enabling robots to autonomously generate safe and feasible motions in dynamic environments remains one of the central challenges in robotic learning and motion planning. Recent advances in deep generative modeling, particularly in diffusion models, have significantly expanded the capabilities of robots to learn diverse and high-dimensional behaviors from demonstrations or interactions. These models have proven effective in modeling multimodal trajectory distributions, enabling robots to adaptively generate context-dependent motions across various tasks, such as navigation [1], object manipulation [2], and locomotion [3].
@@
Despite their success, diffusion models come with notable limitations. The inference process is typically slow and computationally intensive, as it involves iteratively solving a stochastic differential equation (SDE) over many denoising steps. This high inference cost impedes real-time decision making, especially in safety-critical applications where rapid reactions to environmental changes are necessary. Furthermore, while diffusion models can be conditioned on goals or task descriptions, they generally lack formal safety guarantees, which is crucial when deploying robots in dynamic environments with obstacles, humans, or other agents. Recent works have begun to address these concerns by incorporating safety constraints into diffusion-based frameworks (e.g., via model-predictive control or barrier functions), but doing so often further increases computational complexity and may compromise sample quality or task performance.
@@
To overcome these challenges, we turn to Flow Matching (FM)—a recently emerging generative modeling framework that offers a promising alternative to diffusion models. Instead of relying on stochastic processes, FM defines a deterministic flow field—governed by an ordinary differential equation (ODE)—that smoothly transforms samples from a simple prior distribution into target data samples. FM has demonstrated strong performance across several domains, including image synthesis, video generation, speech, proteins, and robot trajectory learning, largely due to its numerical stability, training simplicity, and efficient inference. More importantly, the FM framework is amenable to explicit constraint integration, making it a natural candidate for generating safe robot motions.
@@
In this work, we introduce Safe Flow Matching (SafeFM), a novel method for constraint-aware motion planning that combines the efficiency and flexibility of flow matching with the formal safety guarantees provided by Control Barrier Functions (CBFs). SafeFM learns a continuous-time vector field that maps random waypoints from a prior distribution to goal-directed robot trajectories, while simultaneously ensuring that the generated trajectories respect predefined safety constraints throughout the entire flow. By leveraging finite-time flow invariance principles, we embed safety specifications directly into the learned dynamics, enabling real-time generation of safe and feasible motions without sacrificing generation quality or expressiveness.
To evaluate the efficacy of SafeFM, we conduct extensive experiments across several simulated and real-world scenarios, including planar robot navigation, 7-DoF robotic arm manipulation, and environments with unseen obstacle configurations. We compare SafeFM to both diffusion-based approaches and unconstrained flow-based baselines, demonstrating that our method consistently achieves higher success rates, lower violation of safety constraints, and significantly faster inference times. The main contributions of this paper are:
@@
  • We propose SafeFM, a novel flow matching framework for robot motion planning under safety constraints.
@@
  • We introduce a theoretically grounded integration of control barrier functions into flow-based generative models to ensure safe trajectory generation.
  • We empirically demonstrate the superior safety, efficiency, and generalization capabilities of SafeFM across multiple tasks.
@@ 2 Related Work
Diffusion models have been effectively utilized in various robot learning domains, including motion planning, imitation learning, and reinforcement learning, as demonstrated by works such as [4, 2, 5]. In imitation learning, diffusion models have been leveraged to produce robot trajectories based on expert demonstrations. For example, Pearce et al. [6] and Reuss et al. [5] employ diffusion models to develop policies that replicate expert behaviors. Additionally, Xian et al. [7] introduce ChainedDiffuser, a novel policy architecture that integrates action keypose prediction with trajectory diffusion generation for robotic manipulation tasks. To tackle long-horizon planning and task execution, Li et al. [8] and Mishra et al. [9] propose skill-centric diffusion models that combine learned distributions to create extended plans during inference. For generating control actions, Chi et al. [10] present Diffusion Policy, which frames policy learning as a conditional denoising diffusion process over the robot’s action space, conditioned on 2D observation features. This concept is advanced by Ze et al. [11] in their 3D Diffusion Policy, which incorporates complex 3D visual representations to enable generalizable visuomotor policies from just a few dozen expert demonstrations.
Recent efforts have also aimed to ensure that generated trajectories comply with the physical laws governing robot dynamics. For instance, Gadginmath and Pasqualetti [3] embed system dynamics directly into the denoising process, while Chen et al. [12] use generated state sequences with an inverse dynamics model to derive actions for coordinated bimanual tasks. To enforce safety and dynamical constraints, Li et al. [13] incorporate penalties for constraint violations into the loss function, though this method does not guarantee strict adherence to hard constraints. Alternatively, Bouvier et al. [14] project generated policies onto a dynamically admissible manifold during training and inference. However, this technique underapproximates the reachable set using a polytope, relying on the assumption of smooth nonlinear dynamics and environments. Drawing inspiration from control barrier functions, Xiao et al. [15] propose Safe Diffuser to enforce safety constraints within diffusion models, influencing our focus on safety in robot motion generation. Furthermore, Mizuta and Leung [16] integrate both system dynamics and control barrier functions into the diffusion framework.
Despite these advances, diffusion models struggle in high-dimensional control tasks, such as robotic manipulation, where achieving real-time inference with high-quality actions remains difficult. This challenge stems from the inherent uncertainty and iterative nature of the diffusion process.
@@
Flow matching models have emerged as a promising alternative to diffusion models, offering potential improvements in efficiency and performance for robotic applications. For instance, Zhang and Gienger [17] introduce a flow matching policy that utilizes affordance-based reasoning for manipulation tasks. Departing from image-based inputs, Chisari et al. [18] employ point clouds with conditional flow matching to generate robot actions. To enhance efficiency, Zhang et al. [19] propose a flow policy that uses consistency flow matching with a straight-line flow, enabling one-step generation and striking a balance between speed and effectiveness, though limitations persist in complex manipulation scenarios. Capitalizing on the strengths of flow matching, Braun et al. [20] apply these models to robot states defined on a Riemannian manifold. This approach better captures high-dimensional multimodal distributions, improving the handling of complex robotic systems.
@@ 3.1 Flow Matching Model
A flow matching model is defined through an ODE governed by a time-dependent vector field $\bm{v}^{\bm{\theta}}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ parameterized by a neural network, which induces a time-dependent flow map $\bm{\psi}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, i.e.,
@@
where $\bm{\psi}_{t}(\bm{\tau}):=\bm{\psi}(t,\bm{\tau})$ and $\bm{v}_{t}^{\bm{\theta}}:=\bm{v}(t,\bm{\tau})$ with initial conditions $\bm{\psi}_{0}(\bm{\tau}_{0})=\bm{\tau}_{0}$. We represent a trajectory $\mathcal{T}(t):[0,1]\to\mathbb{R}^{H\times d_{s}}$ as a sequence of states $\mathbf{s}_{t}^{k}\in\mathbb{R}^{d_{s}}$, such that $\mathcal{T}_{t}=[(\mathbf{s}_{t}^{0})^{\top},\dots,(\mathbf{s}_{t}^{H})^{\top}]\in\mathbb{S}^{H+1}\subseteq\mathbb{R}^{(H+1)d_{s}}$, with the dimensionality $d=d_{s}(H+1)$. Considering a start $\mathcal{T}_{0}=\bm{\tau}_{0}$ at $t=0$, the vector field generates the probability path $p_{t}$ if its flow $\bm{\psi}_{t}(\bm{\tau})$ satisfies
$\mathcal{T}_{t}:=\bm{\psi}_{t}(\mathcal{T}_{0})\sim p_{t}~{}\text{for}~{}\mathcal{T}_{0}\sim p_{0}$ (2)
Thus, the velocity field $\bm{v}_{t}^{\bm{\theta}}$ enables sampling from the probability path $p_{t}$ by solving the ODE. The objective of flow matching is to train a parameterized vector field $\bm{v}t^{\bm{\theta}}$ such that its induced flow $\bm{\psi}t$ generates a probability path $p_{t}$, transitioning from an initial distribution $p_{0}=p$ to the target data distribution $p_{1}=q$.
@@ 3.2 Safe Invariance in Dynamical Systems
Consider a control-affine system
@@
where $\bm{x}\in\mathbb{X}\subset\mathbb{R}^{n}$, $\bm{f}(\cdot):\mathbb{X}\to\mathbb{R}^{n}$ and $\bm{g}(\cdot):\mathbb{X}\to\mathbb{R}^{n\times m}$ are locally Lipschitz, and $\bm{u}\in\mathbb{R}^{m}$ is the control input. Moreover, a super level set :=$\mathbb{C}:=\{\bm{x}_{t}\in\mathbb{R}^{n}|h(\bm{x}_{t})\geq 0\}$ is defined via a continuously differentiable equation $h(\cdot):\mathbb{R}^{n}\to\mathbb{R}$, which encode the safety guarantees for Eq. 3.
@@
Given a safe set $\mathbb{C}$, $h(\bm{x})$ is a time-invariant CBF for system Eq. 3 if there exists a class $\mathcal{K}$ function $\alpha(\cdot)$ such that
$\sup_{\bm{u}\in\mathbb{R}^{m}}\{L_{\bm{f}}h(\bm{x})+L_{\bm{g}}h(\bm{x})\bm{u}+\alpha(h(\bm{x}))\}\geq 0$ (4a)
for all $\bm{x}\in\mathbb{C}$, where $L_{\bm{f}}h(\bm{x}),L_{\bm{g}}h(\bm{x})$ denote the Lie derivatives of $h(\bm{x})$ along $\bm{f}(\bm{x})$ and $\bm{g}(\bm{x})$, respectively.
The safety guaranteed by the control barrier function is shown as follows.
Theorem 1 ([21]).
If $h(\cdot)$ is a valid control barrier function and $\bm{x}(0)\in\mathbb{C}$, then $\mathbb{C}$ is forward invariant indicating $\bm{x}(t)\in\mathbb{C}$ for $\forall t\in\mathbb{R}_{0,+}$.
However, the condition of initial safety $\bm{x}_{0}\in\mathbb{C}$ may not be satisfied in some scenarios. Particularly, in robotic motion planning tasks, it is desired that the generated trajectory enters the safe set $\mathbb{C}$ within a prescribed time $T_{d}\in\mathbb{R}_{+}$, and remains in $\mathbb{C}$ afterwards. Thankfully, this task can be defined as prescribed-time safety, i.e.,
Definition 2.
A system in Eq. 3 is prescribed-time safe with a predefined time $T_{d}\in\mathbb{R}_{+}$, if $\bm{x}_{t}\in\mathbb{C}$ for any $t\geq T_{d}$ with arbitrary $\bm{x}_{0}\in\mathbb{X}$.
To ensure safety in flow matching, we draw inspiration from the safety guarantees provided by control barrier functions and propose an adaptation for trajectory generation. The mathematical formulation of this approach is presented in the following section.
@@ 4 Safe Flow Matching
The objective of safe flow matching is to generate a trajectory $\mathcal{T}=[(\mathbf{s}^{0})^{\top},\dots,(\mathbf{s}^{H})^{\top}]^{\top}$ that closely aligns with a target distribution while remaining within a predefined safe set. In this section, we formally define safe flow matching and introduce the concept of flow barrier functions. Furthermore, we propose safety regularization terms designed to enforce compliance with safety constraints during trajectory generation.
@@ 4.1 Safe Flow Matching under Unified Constraints
In this subsection, we consider a safe set $\mathbb{C}_{s}\subseteq\mathbb{S}$ defined for each state $\mathbf{s}^{i}$, $i=0,\dots,H$, characterized by a single continuous function $h(\cdot):\mathbb{S}\to\mathbb{R}$ such that The generated trajectory $\mathcal{T}$ is called as safe, if all generated states $\bm{s}^{i}$ remain in the safe set $\mathbb{C}_{s}$, which is mathematically defined as follows
@@
A trajectory $\mathcal{T}$ with length $H$ is safe w.r.t $h(\cdot)$, if
$h(\bm{E}_{i}\mathcal{T})\geq 0,$ $\forall i\in\{0,1,\cdots,H\}$ (5)
@@
As the trajectory $\mathcal{T}$ is generated via a flow model in finite time, the trajectory safety can be guaranteed. Moreover, unlike conventional CBFs, flow matching must account for potentially unsafe initial trajectories $\mathcal{T}_{0}\sim p_{0}$. Such a safe trajectory generation task in flow matching is defined as a prescribed-time safety problem, where the initial state is required to start in the predefined safe set. Accordingly, we introduce the definition of safe flow-based trajectory generation as follows.
@@
To ensure the safety of the generated trajectory as in Definition 4, we consider the control input $\bm{u}_{t}=[{\bm{u}_{t}^{0}}^{\top},\cdots,{\bm{u}_{t}^{H}}^{\top}]^{\top}$ serving as a regularization term to steer the flow toward a safe manifold. Therefore, the flow model is formulated as
$\frac{\mathrm{d}}{\mathrm{d}t}\bm{\psi}_{t}(\bm{\tau})=\bm{v}_{t}^{\theta}(\bm{\tau})+\bm{u}_{t},~{}\forall t\in[0,1].$ (6)
Drawing inspiration from control barrier functions, we propose a flow matching barrier function (FMBF) to enforce safety constraints in the flow matching framework, which is defined as
@@
A function $h(\cdot)$ is called a flow matching barrier function if it satisfies the following conditions:
@@
    where
    $\varphi(t,h(\bm{s}))=\begin{cases}\varphi_{0},&\text{if~{}}h(\bm{s})\geq 0\\ \varphi_{1}(t),&\text{otherwise}\end{cases}$ (8)
    with $\varphi_{0}\in\mathbb{R}_{+}$ and a blow-up function $\varphi_{1}(\cdot):[0,1)\to\mathbb{R}_{+}$ satisfying $\lim_{t\to 1^{-}}\varphi_{1}(t)=+\infty$.
@@
Build upon the FMBF, the safety guarantee derived in the following theorem.
@@
Suppose $h(\cdot)$ is a valid flow matching barrier function as per Definition 5. Then, the flow matching process is safe according to Definition 4, ensuring that the generated trajectory $\mathcal{T}_{1}$ satisfies the safety condition in Definition 3.
Although valid flow matching barrier functions ensure safety for both the trajectory and the flow matching process, the selection of an appropriate regularization term $\bm{u}_{t}$ for each $t\in[0,1]$ remains a critical challenge. To minimize the influence of $\bm{u}_{t}$ on the distribution’s alignment with the target $p_{1}$, the regularization term should be kept as small as possible in magnitude. To this end, we determine $\bm{u}_{t}$ by solving the following optimization problem:

### v2

@@ Abstract
Recent advances in generative modeling have led to promising results in robot motion planning, particularly through diffusion and flow matching (FM)-based models that capture complex, multimodal trajectory distributions. However, these methods are typically trained offline and remain limited when faced with new environments with constraints, often lacking explicit mechanisms to ensure safety during deployment. In this work, we propose Safe Flow Matching (SafeFlow), a motion planning framework, for trajectory generation that integrates flow matching with safety guarantees. SafeFlow leverages our proposed flow matching barrier functions (FMBF) to ensure the planned trajectories remain within safe regions across the entire planning horizon. Crucially, our approach enables training-free, real-time safety enforcement at test time, eliminating the need for retraining. We evaluate SafeFlow on a diverse set of tasks, including planar robot navigation and 7-DoF manipulation, demonstrating superior safety and planning performance compared to state-of-the-art generative planners. Comprehensive resources are available on the project website: https://safeflowmatching.github.io
@@ 1 Introduction
Figure 1: SafeFlow process illustration.
@@
Enabling robots to autonomously generate safe and feasible motions in dynamic environments remains one of the central challenges in robotic learning and motion planning. Recent advances in deep generative modeling, particularly in diffusion models, have significantly expanded the capabilities of robots to learn diverse and high-dimensional behaviors from demonstrations or interactions. These models have proven effective in modeling multimodal trajectory distributions, enabling robots to adaptively generate context-dependent motions across various tasks, such as navigation [1], object manipulation [2], and locomotion [3]. Despite their success, diffusion models come with notable limitations. Specifically, the inference process is typically slow and computationally intensive, as it involves iteratively solving a stochastic differential equation over many denoising steps [4]. Furthermore, while diffusion models can be conditioned on goals or task descriptions, they generally lack formal safety guarantees, which is crucial when deploying robots in dynamic environments with obstacles, humans, or other agents. Recent works have begun to address these concerns by incorporating safety constraints into diffusion-based frameworks (e.g., via model predictive control [5] or barrier functions [6]), but such repeated inference using diffusion models further increases computational complexity and may compromise inference quality or task performance.
@@
To overcome these challenges, we turn to Flow Matching (FM)—a recently emerging generative modeling framework that offers a promising alternative to diffusion models [7]. Instead of relying on stochastic processes, FM defines a deterministic flow field—governed by an ordinary differential equation—that smoothly transforms samples from a simple prior distribution into target data samples [8]. More importantly, the FM framework is amenable to explicit constraint integration, making it a natural candidate for generating safe robot motions [9, 10] (as shown in Fig. 1).
@@
In this work, we introduce safe flow matching (SafeFlow), a novel method for constraint-aware motion planning that combines the efficiency and flexibility of FM with the formal safety guarantees provided by flow matching barrier functions (FMBFs), inspired by control barrier functions (CBFs). SafeFlow learns a continuous-time vector field that maps random waypoints from a prior distribution to goal-directed robot trajectories, while simultaneously ensuring that the generated trajectories respect safety constraints. By leveraging finite-time flow invariance principles [7], we embed a regularization term directly into the FM model, enabling real-time generation of safe and feasible motions without sacrificing generation quality or expressiveness. We compare SafeFlow to state-of-the-art (SOTA) methods, demonstrating the superiority of our method across six performance metrics, e.g., safety, smoothness, and efficiency. The main contributions of this paper are:
@@
  • SafeFlow, a novel FM framework for robot motion planning under safety constraints.
@@
  • A theoretically grounded integration of control barrier functions, as a plug-and-play safety module, into flow-based generative models to ensure safe trajectory generation.
  • Empirical demonstrations on the superior safety capabilities and planning performance of SafeFlow across multiple tasks.
@@ 2 Related Work
Diffusion models have been effectively utilized in various robot learning domains, including motion planning, imitation learning, and reinforcement learning, as demonstrated by works such as [2, 3, 11]. Especially in imitation learning, diffusion models have been leveraged to produce robot trajectories based on expert demonstrations. For example, Pearce et al. [12] and Reuss et al. [13] employ diffusion models to develop policies that replicate expert behaviors. Additionally, Xian et al. [14] introduces ChainedDiffuser, a novel policy architecture that integrates action keypose prediction with trajectory diffusion generation for robotic manipulation tasks. To tackle long-horizon planning and task execution, Li et al. [15] and Mishra et al. [16] propose skill-centric diffusion models that combine learned distributions to create extended plans during inference. For generating control actions, Chi et al. [17] presents Diffusion Policy, which frames policy learning as a conditional denoising diffusion process over the robot’s action space, conditioned on 2D observation features. This concept is advanced by Ze et al. [18] in their 3D Diffusion Policy, which incorporates complex 3D visual representations to enable generalizable visuomotor policies from just a few dozen expert demonstrations.
Recent efforts have also aimed to ensure that generated trajectories comply with the physical laws governing robot dynamics. For instance, Gadginmath and Pasqualetti [3] embed system dynamics directly into the denoising process, while Chen et al. [19] use generated state sequences with an inverse dynamics model to derive actions for coordinated bimanual tasks. To enforce safety and dynamical constraints, Li et al. [20] incorporate penalties for constraint violations into the loss function, though this method does not guarantee strict adherence to hard constraints. Alternatively, Bouvier et al. [21] project generated policies onto a dynamically admissible manifold during training and inference. However, this technique underapproximates the reachable set using a polytope, relying on the assumption of smooth nonlinear dynamics and environments. Drawing inspiration from CBFs, Xiao et al. [22], Mizuta and Leung [23] propose SafeDiffuser and CoBL, frameworks aimed at enforcing safety constraints within reverse diffusion models for safe planning. However, a critical limitation remains: as acknowledged by the authors, CBFs rely on the assumption that the initial state of the reverse diffusion process lies within the safe set, an assumption that is fundamentally violated, since the initial state is sampled from pure noise. Furthermore, due to the probabilistic nature of diffusion models, there is no guarantee that the generated trajectories will remain within the safe set throughout the process. In addition, the manipulation tasks considered in [22] are overly simplistic, addressing only joint limit constraints while overlooking critical factors such as collision avoidance. As a result, the practical applicability of this framework to real-world robot learning and control remains limited.
@@
Flow matching models have emerged as a promising alternative to diffusion models, offering potential improvements in efficiency and performance for robotic applications. For instance, Zhang and Gienger [24] introduces a flow matching policy that utilizes affordance-based reasoning for manipulation tasks. Departing from image-based inputs, Chisari et al. [25] employs point clouds with conditional flow matching to generate robot actions. To enhance efficiency, Zhang et al. [26] proposes a flow policy that uses consistency flow matching with a straight-line flow, enabling one-step generation and striking a balance between speed and effectiveness. However, limitations persist in complex manipulation scenarios. Capitalizing on the strengths of flow matching, Braun et al. [27] applies these models to robot states defined on a Riemannian manifold. This approach better captures high-dimensional multimodal distributions, improving the handling of complex robotic systems.
@@ 3.1 Flow Matching Model
In this paper, we intend to generate the motion trajectory $\mathcal{T}(\cdot):[0,1]\to\mathbb{R}^{(H+1)\times d_{s}}$ of robotics with a horizon $H\in\mathbb{N}_{>0}$, which is represented as a sequence of states $\mathbf{s}_{t}^{k}\in\mathbb{S}\subseteq\mathbb{R}^{d_{s}}$ for $t\in[0,1]$ and $k=0,\cdots,H$, i.e., $\mathcal{T}_{t}=\mathcal{T}(t)=[(\mathbf{s}_{t}^{0})^{\top},\dots,(\mathbf{s}_{t}^{H})^{\top}]^{\top}\in\mathbb{S}^{H+1}\subseteq\mathbb{R}^{(H+1)d_{s}}$ with the dimension $d_{s}\in\mathbb{N}_{>0}$ and $d=d_{s}(H+1)$. To this end, a flow matching model is defined through an ODE governed by a time-dependent vector field $\bm{v}^{\bm{\theta}}:[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ parameterized by a neural network with parameters $\bm{\theta}$, which induces a time-dependent flow map $\bm{\psi}(\cdot,\cdot):[0,1]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ with $d\in\mathbb{N}_{>0}$ [7]. Specifically, the flow matching model is written as
@@
where $\bm{\psi}_{t}(\bm{\tau}):=\bm{\psi}(t,\bm{\tau})$ and $\bm{v}_{t}^{\bm{\theta}}:=\bm{v}(t,\bm{\tau})$ with initial conditions $\bm{\psi}_{0}(\bm{\tau}_{0})=\bm{\tau}_{0}$. Considering a start $\mathcal{T}_{0}=\bm{\tau}_{0}$ at $t=0$, the vector field generates the probability path $p_{t}$ if its flow $\bm{\psi}_{t}(\bm{\tau})$ satisfies
$\mathcal{T}_{t}:=\bm{\psi}_{t}(\mathcal{T}_{0})\sim p_{t}$ for $\mathcal{T}_{0}\sim p_{0}$ (2)
Thus, the velocity field $\bm{v}_{t}^{\bm{\theta}}$ enables sampling from the probability path $p_{t}$ by solving the ODE. The objective of flow matching is to train a parameterized vector field $\bm{v}_{t}^{\bm{\theta}}$ such that its induced flow $\bm{\psi}_{t}$ generates a probability path $p_{t}$, transitioning from an initial distribution $p_{0}=p$ to the target data distribution $p_{1}=q$.
@@ 3.2 Safe Invariance in Dynamical Systems
To ensure the safety of the generated trajectory $\mathcal{T}$, CBFs are introduced here. In particular, consider a control-affine system
@@
where $\bm{x}_{t}\in\mathbb{X}\subset\mathbb{R}^{n}$, $\bm{f}(\cdot):\mathbb{X}\to\mathbb{R}^{n}$ and $\bm{g}(\cdot):\mathbb{X}\to\mathbb{R}^{n\times m}$ are locally Lipschitz, and $\bm{u}_{t}\in\mathbb{R}^{m}$ is the control input for all time $t$. Moreover, a super level set :=$\mathbb{C}:=\{\bm{x}_{t}\in\mathbb{R}^{n}|h(\bm{x}_{t})\geq 0\}$ is defined via a continuously differentiable equation $h(\cdot):\mathbb{X}\!\to\!\mathbb{R}$, which encodes safety guarantees for Eq. 3.
@@
Given a safe set $\mathbb{C}$, $h(\bm{x})$ is a time-invariant CBF for system Eq. 3 if there exists a class $\mathcal{K}$ function $\alpha(\cdot)$ such that for all $t\in\mathbb{R}_{0,+}$ it has $\bm{x}_{0}\in\mathbb{C}$ and
$\sup_{\bm{u}\in\mathbb{R}^{m}}\{L_{\bm{f}}h(\bm{x}_{t})+L_{\bm{g}}h(\bm{x}_{t})\bm{u}+\alpha(h(\bm{x}_{t}))\}\geq 0,$ (4)
where $L_{\bm{f}}h(\bm{x}_{t}),L_{\bm{g}}h(\bm{x}_{t})$ denote the Lie derivatives of $h(\bm{x}_{t})$ along $\bm{f}(\bm{x}_{t})$ and $\bm{g}(\bm{x}_{t})$, respectively.
Notably, the initial safety condition $\bm{x}_{0}\in\mathbb{C}$ is generally not satisfied in the reverse diffusion process, which has been overlooked in previous works [22, 23]. Thankfully, the task of driving the system Eq. 3 into the safe set during the reverse diffusion process can be defined as prescribed-time safety as follows.
Definition 2 ([28]).
A system in Eq. 3 is prescribed-time safe with a predefined time $T_{d}\in\mathbb{R}_{>0}$, if $\bm{x}_{t}\in\mathbb{C}$ for any $t\geq T_{d}$ with arbitrary $\bm{x}_{0}\in\mathbb{X}$.
To ensure safety in FM, we draw inspiration from the safety guarantees provided by CBFs and propose an adaptation for trajectory generation. The mathematical formulation of this approach is presented in the following section.
@@ 4 Safe Flow Matching
The objective is to generate a trajectory $\mathcal{T}=[(\mathbf{s}^{0})^{\top},\dots,(\mathbf{s}^{H})^{\top}]^{\top}$ that closely aligns with a target distribution while remaining within a predefined safe set. In this section, we formally define safe flow matching and introduce the concept of flow matching barrier functions (FMBFs). Furthermore, we propose safety regularization terms designed to enforce compliance with safety constraints during trajectory generation.
@@ 4.1 Safe Flow Matching under Unified Constraints
In this subsection, we consider a safe set $\mathbb{C}_{s}\subseteq\mathbb{S}$ defined for each state $\mathbf{s}^{i}$, $i=0,\dots,H$, characterized by a single continuous function $h(\cdot):\mathbb{S}\to\mathbb{R}$. Then the generated trajectory $\mathcal{T}$ is called safe if all generated states $\bm{s}^{i}$ remain in the safe set $\mathbb{C}_{s}$, which is mathematically defined as follows.
@@
A trajectory $\mathcal{T}$ with length $H+1$ is safe w.r.t $h(\cdot)$, if
$h(\bm{E}_{i}\mathcal{T})\geq 0,$ $\forall i\in\{0,1,\cdots,H\},$ (5)
@@
Accordingly, we introduce the definition of safe flow matching trajectory generation as follows.
@@
To ensure the safety of the generated trajectory as in Definition 4, we consider the control input $\bm{u}_{t}=[{\bm{u}_{t}^{0}}^{\top},\cdots,{\bm{u}_{t}^{H}}^{\top}]^{\top}$ serving as a regularization term to steer the flow toward a safe manifold, where $\bm{u}_{t}^{i}\in\mathbb{R}^{d_{s}}$ with $i=1,\cdots,H$. Therefore, the flow model is formulated as
$\frac{\mathrm{d}}{\mathrm{d}t}\bm{\psi}_{t}(\bm{\tau})=\bm{v}_{t}^{\theta}(\bm{\tau})+\bm{u}_{t},$ $\forall t\in[0,1].$ (6)
Drawing inspiration from CBFs, we propose an FMBF to enforce safety constraints in the flow matching framework, which is defined as follows.
@@
A function $h(\cdot)$ is called a flow matching barrier function (FMBF) if it satisfies the following conditions:
@@
    with $\varphi(t,h(\bm{s}))=\varphi_{0}\in\mathbb{R}_{>0}$ if $h(\bm{s})\geq 0$, and $\varphi(t,h(\bm{s}))=\varphi_{1}(t)$, otherwise, where $\varphi_{1}(\cdot):[0,1)\to\mathbb{R}_{>0}$ monotonically increases satisfying $\lim_{t\to 1^{-}}\varphi_{1}(t)\!=\!\infty$ and $\int_{0}^{1^{-}}\varphi_{1}(s)\mathrm{d}s\!=\!\infty$.
@@
Build upon the FMBF, the safety guarantee is derived in the following theorem [^1] For the detailed proofs, refer to Appendix A..
@@
Suppose $h(\cdot)$ is a valid FMBF as per Definition 5. Then, the FM process is safe according to Definition 4, ensuring that the generated trajectory $\mathcal{T}_{1}$ satisfies the safety condition in Definition 3.
Although valid FMBFs ensure safety for both the trajectory and the FM process, the selection of an appropriate regularization term $\bm{u}_{t}$ for each $t\in[0,1]$ remains a critical challenge. To minimize the influence of $\bm{u}_{t}$ on the distribution’s alignment with the target $p_{1}$, the regularization term should be kept as small as possible in magnitude. To this end, we determine $\bm{u}_{t}$ by solving the following optimization problem: